OpenStudy (anonymous):

HELP ME PLZ Use the discriminant to determine how many real solutions there are to the function y = -3x2 - 3x - 1. A.There are no real number solutions. B. All real numbers are solutions. C.There are two real number solutions. D.There is only one real number solution. Use the discriminant to determine how many real solutions there are to the function y = -5x2 - 10x - 5. A.There are no real number solutions. B.All real numbers are solutions. C.There are two real number solutions. D.There is only one real number solution. Use the quadratic formula to determine the solutions for the function y = -2x2 + 3x + 2 A.There are no real number solutions. B.x = -1/2, x = 2 C.x = -1/2 D.x = 1/2, x = -2 Use the quadratic formula to determine the solutions for the function y = 3x2 - 5x - 2. A.There are no real number solutions. B.x = 1/3, x = 2 C.x = -1/3 D. x = -1/3, x = 2

OpenStudy (gorv):

if D > 0 then two real solutio D=0 one real solution as both root will be equal D<0 no real solution

OpenStudy (gorv):

ax^2+bx+c=0 D=b^2-4*a*c for Q 1 b=-3 a=-3 c=-1 D=(-3)^2-4*(-1)*(-3)=9-12=-3<0 so no real solution